if 49.0 g of H3PO4 is reacted with excess KOH, determine the theoretical yield of K3PO4 if you isolate 49.0g of K3PO4?
1
Expert's answer
2021-10-19T13:28:03-0400

H3PO4 + 3KOH → K3PO4 + 3H2O

M(H3PO4) = 98.0 g/mol

n(H3PO4) "= \frac{49.0}{98.0}= 0.5 \;mol"

According to the reaction:

n(H3PO4)=n(K3PO4) = 0.5 mol

M(K3PO4) = 212.0 g/mol

m(K3PO4) "= 0.5 \times 212.0 = 106.0 \;g"

Proportion:

106.0 g – 100 %

49.0 g – x

"x = \frac{49.0 \times 100}{106.0} = 46.22 \; \%"

Answer: 46.22 %